# The Curious Case of Arbitrariness in Machine Learning

### v1

@@ 6.2. Multiplicity and Homogenization
Term Definition
Algorithmic Blackballing (Ajunwa, 2020) ‘A worker’s lack of control over the portability of applicant data captured by automated hiring systems […] raising the specter of an algorithmically permanently excluded class of job applicants’
Algorithmic Monoculture (Kleinberg and Raghavan, 2021) ‘The notion that choices and preferences will become homogeneous in the face of algorithmic curation’
Algorithmic Leviathan (Creel and Hellman, 2022) ‘Automated decision-making systems that make uniform judgments across broad swathes of a sector.’
Outcome Homogenization (Bommasani et al., 2022) ‘The phenomenon of individuals (or groups) exclusively receiving negative outcomes from all ML models they interact with’
Generative Monoculture (Wu et al., 2024) ‘A distribution shift [towards less varied outputs] from source data (i.e., human-generated training data) to model-generated data (i.e., model outputs) for a specific task.’
Algorithmic Pluralism (Jain et al., 2024b) ‘A state of affairs in which the algorithms used for decision-making are not so pervasive and/or strict as to constitute a severe bottleneck on opportunity.’
Table 2. Various terms used in the homogenization literature.
Interestingly, introducing controlled arbitrariness or multiplicity can mitigate these risks by preventing monocultures. In contexts where arbitrariness in individual decisions is less concerning but homogenization is, controlled multiplicity becomes a desirable property (Barocas et al., 2023; Perry and Zarsky, 2015; Jain et al., 2024a). In such situations, when an intentional design choice is not possible, the developers should prefer arbitrary choices over conventional ones wherever feasible (§2.1). Despite this idea being widely recognized in the academic literature (Barocas et al., 2023; Perry and Zarsky, 2015; Jain et al., 2024a), public perception of intentional randomness in decision-making remains skeptical. A recent study by Meyer et al. (2024) indicates a strong aversion of the end users towards any form of randomization or intentional arbitrariness in automated decision-making. Therefore, fostering greater public awareness about the nuanced impacts of multiplicity is crucial before we can develop and employ potential solutions (Barocas et al., 2023).
7. Open Research Questions and Emerging Trends
Building upon our systematic survey of the literature on multiplicity, this work gives us a unique vantage point to identify and discuss several emerging trends and critical open research questions in the field,
  • Expanding the scope of multiplicity beyond predictions and explanations. One of our key motivations for broadening the definition of multiplicity (§3.1) was to incorporate multiplicity beyond predictive and explanatory contexts. Although interest in these aspects is growing, further work is needed to bring them to the forefront. We hope our work will encourage future research that fosters a deeper and more holistic understanding of multiplicity’s broader impact.
@@ 7. Open Research Questions and Emerging Trends
  • Mathematical foundations of multiplicity. Establishing stronger mathematical foundations of multiplicity, for instance, our focus on formalizing the distinction between multiplicity, uncertainty, and bias-variance decomposition (§4), is essential. Fundamental work on the Rashomon effect in machine learning is less represented (only $12.5\%$; see Figure 1), highlighting the opportunities for future work on frameworks that rigorously define and explore multiplicity.
@@
  • Interdisciplinary perspective on multiplicity. While substantial work has been done on measuring and mitigating multiplicity, many studies do not engage further with the concerns of multiplicity in real-world settings. As discussed, arbitrariness can be both a force of good (§6.2) or a cause of harm (§6.1), depending on the context. Our systematic review shows that only $47.5\%$ of works explicitly engage with responsible AI concerns (see Figure 1), which we believe to be low given the field’s relevance to the responsible deployment of machine learning models. Future collaborative efforts across disciplines are crucial to truly grasp the real-world consequences of multiplicity.
  • Multiplicity in the era of LLMs. As machine learning models continue to scale, particularly in the era of large language models (LLMs), new challenges emerge. For instance, we see increasing attention given to the concerns of monoculture and homogenization (§6.2). Even the evaluation of multiplicity becomes increasingly complex, as training multiple models is often infeasible at this scale. Additionally, we see another dimension of multiplicity in LLMs–decisions multiplicity due to changing prompts–that requires deeper examination (Lu et al., 2022; Alzahrani et al., 2024; Sclar et al., 2023; Voronov et al., 2024).
8. Conclusion
In this work, we systemized the existing knowledge on multiplicity, uncovering interesting trends and insights. One limitation of our study is the evolving terminology within the field–terms such as “multiplicity” and “Rashomon sets” have gained prominence only recently. As a result, our survey may have missed relevant works that did not explicitly use this terminology. Despite this, our efforts to formalize key discussions–the language around developer choices, definitions, and the distinction between multiplicity, uncertainty, and variance–represent a crucial step toward unifying the field. We also explored broad trends related to the real-world impacts of multiplicity, building from existing literature and highlighting overarching themes that extend beyond their originally studied contexts. We hope our work provides a platform that is both accessible to newcomers and valuable to experts, fostering further research in multiplicity.
Acknowledgements.
Funding support for project activities has been partially provided by Canada CIFAR AI Chair, Google award, FRQNT and NSERC Discovery Grants program. We also express our gratitude to Compute Canada for their support in providing facilities for our evaluations.
@@ A.1. Paper Collection
Table 3. Paper collection statistics.
A.2. Manual Filtering
@@
Table 4. Inclusion and exclusion criteria for manual filtering.
A.3. Manual Tagging
@@
Table 5. Rules for each tag in the manual tagging step.

### v2

@@ 6.2 Multiplicity and Homogenization
Introducing controlled randomness can mitigate these risks by preventing monocultures. In contexts where arbitrariness in individual decisions is less concerning than homogenization, controlled multiplicity is desirable (Barocas, Hardt, and Narayanan 2023; Perry and Zarsky 2015; Jain, Creel, and Wilson 2024). In such situations, when an intentional design choice is not possible, the developers should prefer arbitrary choices over conventional ones wherever feasible (§2.1). Despite this idea being widely recognized in the academic literature (Barocas, Hardt, and Narayanan 2023; Perry and Zarsky 2015; Jain, Creel, and Wilson 2024), public perception of intentional randomness in decision-making remains skeptical. A recent study by Meyer et al. (2024) indicates a strong aversion of the end users towards any form of randomization or intentional arbitrariness in automated decision-making. Therefore, fostering greater public awareness about the nuanced impacts of multiplicity is crucial before we can develop and employ potential solutions (Barocas, Hardt, and Narayanan 2023).
7 Open Questions and Emerging Trends
Our systematic survey gives us a unique vantage point to identify and discuss several emerging trends in the field.
  • Expanding the scope of multiplicity beyond predictions and explanations. Our key motivation for broadening the definition of multiplicity (§3.1) was to incorporate multiplicity beyond predictive and explanatory contexts. Although interest in these aspects is growing, further work is needed to bring them to the forefront. We hope our work will encourage future research that fosters a more holistic understanding of multiplicity’s broader impact.
@@ 7 Open Questions and Emerging Trends
  • Mathematical foundations of multiplicity. Establishing stronger mathematical foundations of multiplicity, for instance, our focus on formalizing the distinction between multiplicity, uncertainty, and bias-variance decomposition (§4), is essential. Fundamental work on the Rashomon effect is less represented (only $12.5\%$; see Figure 1), highlighting the opportunities for future work on frameworks that rigorously define and explore multiplicity.
@@
  • Interdisciplinary perspective on multiplicity. We found many studies in this field do not engage with the concerns of multiplicity in real-world settings. Our systematic review shows that only $47.5\%$ of works explicitly engage with responsible AI concerns (see Figure 1), which we believe to be low given the field’s relevance to the responsible deployment of machine learning models. Future collaborative efforts across disciplines are thus crucial.
  • Multiplicity and LLMs. As models continue to scale, new challenges emerge. For instance, we see increasing attention given to the concerns of monoculture and homogenization (§6.2). Even the evaluation of multiplicity becomes increasingly complex, as training multiple models is often infeasible at this scale. Additionally, we see new dimensions of multiplicity like prompt multiplicity (Ganesh, Shokri, and Farnadi 2025), preference multiplicity, etc., that require deeper examination.
8 Conclusion
In this work, we systemized existing knowledge on multiplicity, uncovering interesting trends. One limitation of our study is the evolving terminology within the field–terms such as “multiplicity” and “Rashomon sets” have gained prominence only recently. As a result, our survey may have missed relevant works that did not explicitly use this terminology. Despite this, our efforts to formalize key discussions–the language around developer choices, definitions, and the distinction between multiplicity, uncertainty, and variance–represent a crucial step toward unifying the field. We also explored broad trends related to the real-world impacts of multiplicity, building from existing literature and highlighting overarching themes that extend beyond their originally studied contexts. We hope our work provides a platform that is both accessible to newcomers and valuable to experts, fostering further research in multiplicity.
Acknowledgements
We would like to thank Shomik Jain for his feedback on an earlier version of the paper. Funding support for project activities has been partially provided by Canada CIFAR AI Chair, Google award, CIFAR Catalyst Grant award, FRQNT and NSERC Discovery Grants program. We also express our gratitude to Compute Canada for their support in providing facilities for our evaluations.
@@ Appendix A Systematic Literature Review
Table 3: Paper collection statistics.
@@
Table 4: Inclusion and exclusion criteria for manual filtering.
  1. 1.
  2. 2.
  3. 3.
A.1 Paper Collection
A.2 Manual Filtering
A.3 Manual Tagging
@@
Table 5: Rules for each tag in the manual tagging step.
Appendix B Metric Glossary
  • Ambiguity (Marx, Calmon, and Ustun 2020)
  • Obscurity (Cavus and Biecek 2024a)
  • Discrepancy (Marx, Calmon, and Ustun 2020)
  • Degree of Underspecification (Teney, Peyrard, and Abbasnejad 2022)
  • Viable Prediction Range (Watson-Daniels, Parkes, and Ustun 2023)
  • Rashomon Capacity (Hsu and Calmon 2022)
  • Multi-target Ambiguity (Watson-Daniels et al. 2023)
  • Rank List Sensitivity (Oh et al. 2022)
  • Std. of Scores (Heljakka et al. 2022; Long et al. 2024)
  • Representational Multiplicity (Heljakka et al. 2022)
  • Region Similarity Score (Somepalli et al. 2022)
  • Self-consistency (Cooper et al. 2024)
  • $\epsilon$-robust to Dataset Multiplicity (Meyer, Albarghouthi, and D’Antoni 2023)
  • Unfairness Range (Mata, Kanamori, and Arimura 2022)
  • Rashomon Ratio (Semenova, Rudin, and Parr 2022)
  • Underspecification Score (Madras, Atwood, and D’Amour 2019)
  • Accuracy Under Intervention (Ganesh 2024)
  • Consistency (Watson, Hasan, and Al Moubayed 2022)
  • Model Class Reliance (Fisher, Rudin, and Dominici 2019)
  • Attribution Agreement (Krishna et al. 2022; Brunet, Anderson, and Zemel 2022)
  • Profile Disparity Index (Kobylińska et al. 2024)
  • Inv. Cost of Neg. Surprise (Pawelczyk, Broelemann, and Kasneci 2020)
  • Variable Importance Clouds (Dong and Rudin 2019)
  • Coverage & Interval Width (Marx et al. 2023)